what are the measures of two supplementary angles M and N if M=3x-6 and N=5x+2
Answer by user_dude2008(1825)

3x-6+5x+2=180
8x-4=180
8x=180+4
8x=184
x=23
M=3(23)-6=69-6=63
N=5(23)+2=115+2=117